The following data is obtained from the general records in the shipping department at United Federal Delivery Company for August:

Inspecting packages prior to placement on truck takes 1 minute per package. Processing paperwork for each shipment (each truck loaded) takes 30 minutes.

Loading packages takes 1 minute of labor per package plus an additional 3 labor hours per truck

Miscellaneous tasks take 2 hours per truck.

Temporary employees to perform all of the above tasks can be hired for any amount of hours at a rate of $20 per hour during the summer (no benefits).

One supervisor is needed for every 1,000 hours of temporary employee labor (rounded up).

Supervisors are paid a flat rate of $4,100 per month.

Supervisors divide their time evenly among inspecting, processing, and loading activities, and miscellaneous tasks.

$1,000 per month general department overhead is allocated to miscellaneous tasks and $2,000 per month to loading, inspecting and processing activities.

August will have 12 working days. United Federal expects to load 6 trucks per day during August. There are on average 100 packages per truck.

Required:


Prepare a budget using the activity-based approach for the costs of inspecting, processing, loading,

and miscellaneous activities within the shipping department for August.
Answer #1

Calculation of no.of hours worked in the month of August:-

Note:- Assuming temporary labours are employed for the remaining days of the month i.e., 30-12 = 18 days.

Particulars No.of hours worked per day No.of hours worked per month

Inspection 1*100*6 = 600min = 10hrs 10* 18 = 180hrs

Processing 30min*6 = 180min = 3hrs 3*18 = 54hrs

Loading (1*100min*6) +(3*6hrs) = 10hrs+18hrs = 28hrs 28*18 = 504hrs

Miscelleneous 2hrs*6 = 12hrs 12*18 = 216hrs

Total no.of hours worked in the month of August = 180+54+504+216 = 954hrs = 1000hrs(rounded up)

  BUDGET USING ACTIVITY BASED APPROACH:-

Particulars Labour cost($) Overhead cost($) Supervisor cost($) Total cost($)

Inspection 3600 2000 4100/4 = 1025 6625

Processing 1080 2000 1025 4105

Loading 10080 2000 1025 13105

Miscelleneous 4320 1000 1025 6345

30180